OpenStudy (anonymous):

How do you solve 5x^2 = 2x +1 by completing the square?

OpenStudy (bahrom7893):

firs move everything to the left: 5x^2 - 2x -1 = 0

OpenStudy (anonymous):

okay

OpenStudy (bahrom7893):

divide everything by 5

OpenStudy (anonymous):

2x^2-5x-3 = 0 (2x + 1)(x - 3) = 0 x = -1/2 or x = 3

OpenStudy (bahrom7893):

x^2 - (2/5)x - 1/5 = 0

OpenStudy (bahrom7893):

So now u have the form of x^2 - 2ax +a^2, and ur a = 1/5

OpenStudy (anonymous):

i already have the answer i just cant seem to get the answer

OpenStudy (bahrom7893):

5x^2 - 2x - a^2 +a^2 - 1 = 0 5x^2 - 2x - 1/25 + 1/25 - 1 = 0

OpenStudy (bahrom7893):

rewrite: sorry the 5s in front of x shouldnt be there

OpenStudy (bahrom7893):

x^2 - 2x - a^2 +a^2 - 1 = 0 x^2 - 2x - 1/25 + 1/25 - 1 = 0 x^2 - 2x + 1/25 - 1/25 - 1 = 0

OpenStudy (anonymous):

?????????

OpenStudy (anonymous):

okay now what

OpenStudy (anonymous):

i got some thing diff whow did you get that

OpenStudy (bahrom7893):

what the heck, now i forgot the 5s, ughh sorry lets start over..

OpenStudy (anonymous):

ahaa would it help if you had the answers

OpenStudy (bahrom7893):

no, i know them, i just hate typin out math solutions

OpenStudy (anonymous):

it says in the book that the answer is x+ 1/5 plus or minus the square root od 5/2

OpenStudy (anonymous):

please :)

OpenStudy (bahrom7893):

5x^2 - 2x -1 = 0 x^2 - (2/5)x + 1/5 = 0 x^2 - (2/5)x + 1/25 - 1/25 + 1/5 = 0

OpenStudy (anonymous):

2x^2-5x-3 = 0 (2x + 1)(x - 3) = 0 x = -1/2 or x = 3

OpenStudy (bahrom7893):

x^2 - 2 (1/5) x + (1/5)^2 - 1/25 + 1/5

OpenStudy (anonymous):

that is what i got but i can be wrong

OpenStudy (bahrom7893):

ur copying the wrong question dkglenn, stop confusing us

OpenStudy (anonymous):

dkglenn ur wrong sorry

OpenStudy (anonymous):

oh ok

OpenStudy (bahrom7893):

so: x^2 - 2 (1/5) x + (1/5)^2 - 1/25 + 1/5 = 0

OpenStudy (bahrom7893):

see u have to recognize that: x^2 - 2 (1/5) x + (1/5)^2 is the same as: (x - 1/5)^2

OpenStudy (anonymous):

ok

OpenStudy (bahrom7893):

(x - 1/5)^2 - 1/25 + 1/5 = 0 (x - 1/5)^2 = 1/25 - 1/5

OpenStudy (bahrom7893):

wait smth went wrong lemme check my arithmetic

OpenStudy (anonymous):

mkayy

OpenStudy (anonymous):

ive done the problem 5 diff times..so stressed so thanks for trying

OpenStudy (bahrom7893):

okay so let me post the problem over again, thread is too long

OpenStudy (anonymous):

mkay